10.07

The Evanec Companys next expected dividend, D1, is $2.70; its growth rate is 4%; and its common stock now sells for $40.00.
Answer #1

Cost of retained earnings=(D1/Current price)+Growth rate

=(2.7/40)+0.04

=10.75%

% flotation cost=(Current price-Net price)/Current price

=(40-38)/40

=5%

Cost of new common stock=(D1/Net price)+Growth rate

=(2.7/38)+0.04

=11.11%(Approx)